# RankByGene: Gene-Guided Histopathology Representation Learning Through Cross-Modal Ranking Consistency

### v1

@@ 4 Experiments and Results
4.1 Dataset
@@
In this study, we use the breast and lung ST data from the HEST-1k dataset [30], which collects publicly available, high-quality ST datasets and applies standardized processing to all collected data. For WSIs, we utilize the TCGA-BRCA (Breast Invasive Carcinoma) and TCGA-LUAD (Lung Adenocarcinoma) datasets from TCGA cohorts [1]. Detailed descriptions of each dataset will be provided in the following sections.
Breast ST datasets. The Breast ST dataset used for training [7] contains 36 ST samples. Each sample contains approximately 15,000 genes, and after preprocessing, we obtained 13,620 spots from all ST samples. Following the external evaluation approach described in [17], we selected two Visium breast ST samples from 10x Genomics to evaluate our model’s performance in predicting gene expression. The first sample is the Human Breast Cancer: Visium Fresh Frozen, Whole Transcriptome [3], containing 4,898 spots. The second sample is the Human Breast Cancer (Block A Section 1) Visium [2], containing 3,813 spots. We refer to the first sample [3] as Breast-ST1 and the second sample [2] as Breast-ST2.
Lung ST datasets. The Lung ST dataset used for training originates from [47] and consists of 6 ST samples, each containing approximately 18, 000 genes, with a total of 12,891 spots. For testing, we select 4 lung ST samples from [67], which we refer to as Lung-ST, containing a total of 1,831 spots.
Cancer-Specific Gene List Selection To manage the high dimensionality of gene expression data, existing methods typically select a list of highly expressed genes for gene-image alignment, such as choosing the top 250 [17], top 100 [12], or top 50 [73, 30, 17] genes. However, this approach may overlook low-expression but significant genes, which can be crucial for downstream WSI analysis tasks. To help the model align images with cancer-specific genes, we select prognosis-related genes from the Human Protein Atlas [4]. Specifically, we first selected genes related to breast and lung cancer and then intersected these with genes in the ST data, resulting in final gene lists of 447 genes for breast ST and 1916 genes for lung ST data. We utilized these gene lists to train both our model and all multi-modal baseline methods. We will provide more details about the gene lists in the supplementary materials.
WSI Datasets. We use survival prediction on TCGA-BRCA (1, 022 WSIs) and TCGA-LUAD (515 WSIs) as our downstream task. Following the preprocessing steps in [42], we cut patches at $224\times 224$ pixels with 20×20$\times$20 × magnification.
4.2 Evaluation Metrics
@@
For the gene expression prediction task, following previous methods [73, 17, 30], we use Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) as evaluation metrics to evaluate. For the survival prediction task, we adopt a 5-fold cross-validation setting as in [13, 76, 81] and use the cross-validated Concordance Index (C-Index) as the evaluation metric.
4.3 Implementation Details
A typical pipeline for our method consists of two stages. In the training stage, we first filter the genes using the selected cancer-related gene list, followed by multi-modal training with Breast ST [7] and Lung ST [47] dataset. In the testing stage, the trained image encoder extracts features for downstream tasks. The image encoder trained on Breast ST [7] dataset is evaluated on Breast-ST1, Breast-ST2, and TCGA-BRCA datasets, while the image encoder trained on Lung ST [47] dataset is evaluated on Lung-ST and TCGA-LUAD datasets.
Image Encoder Breast-ST1 Breast-ST2 Lung-ST
ResNet-50 [23] 0.3633 0.2716 0.0867 0.4042 0.2982 0.1265 0.3105 0.2043 0.0345
CTransPath [70] 0.3572 0.2698 0.0983 0.3994 0.2957 0.1423 0.2945 0.1911 0.0395
UNI [15] 0.3559 0.2648 0.1063 0.4096 0.3178 0.1398 0.2731 0.1724 0.0432
HEST-FT [30] 0.3469 0.2552 0.1269 0.4071 0.3079 0.1592 0.2926 0.1874 0.0682
BLEEP [73] 0.3475 0.2571 0.1198 0.3982 0.2954 0.1437 0.2879 0.1806 0.0751
RankByGene 0.3412 0.2487 0.1803 0.3964 0.2927 0.1762 0.2662 0.1677 0.1019
@@ 4.3 Implementation Details
ResNet-50 [23] 0.5859 $\pm$ 0.0521 0.5631 $\pm$ 0.0443
CTransPath [70] 0.6413 $\pm$ 0.0469 0.5742 $\pm$ 0.0363
UNI [15] 0.6684 $\pm$ 0.0432 0.5865 $\pm$ 0.0254
HEST-FT [30] 0.6531 $\pm$ 0.0453 0.5728 $\pm$ 0.0321
BLEEP [73] 0.6716 $\pm$ 0.0394 0.5778 $\pm$ 0.0383
RankByGene 0.6814 $\pm$ 0.0512 0.5945 $\pm$ 0.0293
Table 1: Results for the tasks of gene expression prediction and survival analysis using different patch encoders. We use Breast-ST1, Breast-ST2, and Lung-ST datasets for gene expression prediction. The slide-level prediction is performed using the ABMIL [28] framework.
We use UNI [15] as the backbone for the image encoder, followed by a 3-layer MLP as the projection head. A 3-layer MLP is also employed for the gene encoder, consistent with the previous method [73]. Both image and gene feature dimensions are set to 1024. For the image encoder, the learning rate is set to 0.0001, while for the gene encoder, it is set to 0.001. The other training settings and test setting are detailed in the supplementary materials.
4.4 Performance Comparison
We compare our method with image-only pretrained baselines, including ResNet-50 [23] (pretrained on ImageNet), CTransPath [70] and UNI [15] (both pretrained on pathology images), as well as gene-guided image feature learning baselines HEST-FT [30] and BLEEP [73].
Gene Prediction. As shown in Figure 4(a), RankByGene outperforms all baseline methods across the Breast-ST1, Breast-ST2, and Lung-ST datasets. Specifically, RankByGene achieves $11\%$ to $42\%$ improvement in PCC over the best baseline performance in each dataset. These results indicate that our method achieves stronger alignment between the image and gene expression data, with the obtained gene-informed image features effectively capturing the underlying global relationships between spots.
Survival Analysis. Using the models trained on Breast ST and Lung ST datasets, we evaluate their performance on survival prediction using TCGA-BRCA and TCGA-LUAD WSIs, respectively. Figure 4(b) presents the performance of different methods on these two datasets. We use the image encoder from each method to extract features on patches, then apply the ABMIL [28] framework to aggregate all patch-level features for slide-level prediction. The table shows that RankByGene achieves the highest C-Index of 0.6814 on the TCGA-BRCA dataset and 0.5945 on the TCGA-LUAD dataset compared to all baseline methods. The strong performance on these two datasets suggests that the improved image-gene alignment enables the model to capture survival-related signals more effectively, highlighting the potential of our method to enhance performance on challenging downstream tasks.
4.5 Ablation Study
@@
ResNet50 0.3633 0.2716 0.0867
ResNet50 + Ours 0.3573 0.2682 0.1035
UNI 0.3559 0.2648 0.1063
UNI + Ours 0.3412 0.2487 0.1803
@@
Table 2: RankByGene ablation experiments. We conduct all ablation experiments on the gene expression prediction task and the Breast-ST1 dataset.
Weight of Cross-Modal Ranking Consistency Loss. We compare the effect of different values for the weight of the ranking consistency loss $\lambda_{1}$ on the Breast-ST dataset. The results in Figure 5(a) show that the model is robust to the choice of the hyperparameter. The performance across all the tested values $\lambda_{1}>0$ significantly outperforms the baselines, with the best result using $\lambda_{1}=5$.
Weight of Intra-Modal Distillation Loss. Figure 5(b) presents the ablation study results for the image consistency loss weight $\lambda_{2}$ on the Breast-ST1 dataset. Our method shows stable performance across MAE, MSE, and PCC at various values of $\lambda_{2}$, with consistent improvements over the baseline methods. These results indicate the benefits of incorporating image consistency loss, as it further enhances the robustness of image feature learning during alignment with gene features.
Ablation Study on Loss Component. Figure 5(c) presents the ablation study results on the impact of individual loss components on the Breast-ST1 dataset. When both loss components are removed, the model exhibits the lowest PCC and higher MAE and MSE values. Including either $\mathcal{L}_{\text{rank}}$ or $\mathcal{L}_{\text{distil}}$ enhances performance, with the best results achieved when both components are applied. The results indicate that both the ranking loss and the image consistency loss contribute significantly to the model’s performance.
Figure 5(d) presents the results of applying our method to different image encoders on the Breast ST1 dataset, specifically ResNet50 (a CNN-based model pretrained on ImageNet) and UNI (a ViT-based model pretrained on pathology images using the DINO framework [11]). The results demonstrate that our method consistently enhances the performance of each encoder in the gene prediction task, as evidenced by lower MAE and MSE values and higher PCC scores. This finding shows the adaptability of our approach, demonstrating its effectiveness across various encoder architectures.
@@ 4.6 Qualitative Result
Figure 5: Visualization of PTGES3 gene expression predictions from different methods, with all values normalized to the range of 0 to 1. Compared to the baseline, our predictions show the closest match to the ground truth.
Rank Accuracy During Training. We define Rank Accuracy as a metric to evaluate whether the ranking relationships among gene features are effectively captured by the image features during training. After the end of each epoch, Rank Accuracy is calculated as follows: the first gene-image pair in each batch is randomly selected as the anchor pair, along with two additional target pairs. Rank Accuracy is determined by checking whether the target pair with higher gene feature similarity to the anchor pair also has a higher image feature similarity. This process is repeated eight times to calculate the final Rank Accuracy. In Figure 6, we compare the change of the rank accuracy during training between different methods. The figure shows that our method achieves a larger improvement in accuracy during training compared to the baseline methods, suggesting that the rank relationships in the gene features are effectively captured by image features during the training process.
Figure 6: Rank accuracy for different methods during training.
Visualization of Cancer Marker Gene. To evaluate the accuracy of our model’s predictions for critical genes, we selected PTGES3 as the marker gene for visualization on the Breast-ST1 dataset due to its association with poor prognosis in breast invasive carcinoma across two independent cohorts [4]. We visualized the gene prediction values for the PTGES3 from different methods across each spot, including ground truth values and the corresponding WSI region, as shown in Figure 5. Compared to the baseline method, our approach provides more accurate predictions for the PTGES3, demonstrating its effectiveness in capturing important gene expression patterns. We also provide visualizations of other marker genes for lung cancer in the supplementary materials.
@@ 5 Conclusion
Our proposed framework establishes a robust and scalable approach for efficiently aligning spatial transcriptomics data with histopathology images, overcoming key challenges in cross-modal alignment. We propose a ranking-based mechanism to capture similarity relationships between gene and image features, facilitating alignment that is resilient to distortions and effective across both local and global scales. This approach preserves essential spatial and molecular patterns, which is critical for accurately capturing the biological complexity of tissue structures. Additionally, our use of a self-supervised knowledge distillation method further enhances the alignment process, enabling the model to adaptively handle the high dimensionality, sparsity, and noise inherent in gene expression data. Through comprehensive experiments, we demonstrate our framework not only improves alignment quality but also enhances predictive capabilities in gene expression prediction and survival analysis. Ultimately, this work provides a foundational tool for multi-modal representation learning in digital pathology.

### v2

@@ 4 Experiments and Results
We evaluate our method on gene expression prediction, and downstream tasks including gene mutation prediction, receptor status classification, and survival analysis.
@@
Datasets and preprocessing. We use the breast and lung ST data from the HEST-1k dataset [32], which collects publicly available, high-quality ST datasets and applies standardized processing to all collected data. For WSIs, we utilize datasets from TCGA cohorts [1] and BCNB [78] for classification and survival prediction tasks. Further details of the datasets are provided in the supplementary.
Following [32], we apply a three-step preprocessing pipeline: L1 normalization, log transformation, and 8-neighborhood smoothing to all ST data. The processed data is used in both training and testing.
Baselines. We compare our method with image-only pretrained baselines, including ResNet-50 [25] (pretrained on ImageNet), CTransPath [74], and UNI [15] (all pretrained on pathology images), as well as gene-guided image feature learning baselines HEST-FT [32] and BLEEP [77]. Since HEST-FT and BLEEP are both gene-image pretraining methods, for a fair comparison we used UNI for the image encoder initialization and use the original initialization for the other baselines. We also compare our method with methods that are specifically designed for gene expression prediction, such as ST-Net [24] and HisToGene [56], using their default initialization.
@@
Evaluation metrics. For the gene expression prediction task, we adopt Mean Squared Error (MSE), Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) as evaluation metrics, following previous studies [77, 17, 32]. The WSI classification performance is assessed using the Area Under the Receiver Operating Characteristic Curve (AUC). For survival prediction, we follow [13, 81, 86] and evaluate the model using the Concordance Index (C-Index).
Image Encoder BCNB TCGA-LUAD mutation
BR PR HER2 EGFR KRAS STK11 TP53
ResNet-50 [25] 0.833 ± 0.039 0.785 ± 0.032 0.709 ± 0.031 0.729 ± 0.068 0.643 ± 0.048 0.739 ± 0.015 0.691 ± 0.023
CTransPath [74] 0.857 ± 0.023 0.792 ± 0.015 0.716 ± 0.019 0.812 ± 0.024 0.682 ± 0.059 0.797 ± 0.025 0.783 ± 0.022
UNI [15] 0.891 ± 0.027 0.809 ± 0.018 0.724 ± 0.032 0.845 ± 0.047 0.712 ± 0.046 0.843 ± 0.021 0.827 ± 0.019
HEST-FT [32] 0.901 ± 0.021 0.811 ± 0.013 0.721 ± 0.015 0.851 ± 0.036 0.716 ± 0.051 0.851 ± 0.028 0.834 ± 0.026
BLEEP [77] 0.899 ± 0.042 0.817 ± 0.019 0.734 ± 0.024 0.862 ± 0.028 0.713 ± 0.049 0.855 ± 0.019 0.838 ± 0.029
RankByGene 0.915 ± 0.024 0.829 ± 0.012 0.759 ± 0.028 0.855 ± 0.037 0.729 ± 0.065 0.861 ± 0.024 0.845 ± 0.033
@@
ResNet-50 [25] 0.586 $\pm$ 0.052 0.563 $\pm$ 0.044
CTransPath [74] 0.641 $\pm$ 0.047 0.574 $\pm$ 0.036
UNI [15] 0.668 $\pm$ 0.043 0.587 $\pm$ 0.025
HEST-FT [32] 0.653 $\pm$ 0.045 0.573 $\pm$ 0.032
BLEEP [77] 0.672 $\pm$ 0.039 0.578 $\pm$ 0.038
RankByGene 0.681 $\pm$ 0.051 0.595 $\pm$ 0.029
Table 2: Classification and survival analysis results on BCNB and TCGA Datasets. Slide-level prediction is performed with ABMIL [30].
@@
ResNet50 0.3812 0.2914 0.0891
ResNet50 + Ours 0.3709 0.2833 0.1123
UNI 0.3692 0.2821 0.1163
UNI + Ours 0.3542 0.2723 0.1874
@@
Table 3: Ablation studies. We conduct all ablation experiments on the gene expression prediction task on Breast-ST1 dataset.
Implementation details. In the training stage, We use UNI [15] as the backbone for the image encoder, followed by a 3-layer MLP as the projection head. A 3-layer MLP is also employed for the gene encoder, the same as the previous method [77]. Following [77], we use cosine similarity to calculate the distance between two modalities.
In the testing stage, we perform 5-fold standalone cross-validation for the gene prediction task. The training set is split into five folds, and the best model from each fold is evaluated on the test set. For classification and survival prediction tasks, we conduct standard 5-fold cross-validation and report both the mean and standard deviation. During inference, we use the teacher model as it performs better. Other training settings, test settings, and resource requirements are provided in the supplementary materials.
Cancer-specific gene selection. Since ST gene expression is high dimensional and extremely sparse, existing methods typically select highly expressed genes for learning and prediction, e.g., top 250 [17], top 100 [12], or top 50 [77, 32, 17] genes. While these highly expressed genes ensures better quality gene features, we are concerned that they may not be relevant to the biology of interest, and thus may be suboptimal for our downstream tasks.
Consulting pathologists, we select prognosis-related genes from the Human Protein Atlas [4] as our cancer-specific gene lists. This resulted in 447 genes for breast cancer and 1916 genes for lung cancer. These gene lists were used for training and evaluation for downstream tasks including gene mutation prediction, receptor status classification, and survival analysis. Despite the concern than some of the genes may be close to zero, we observe these cancer-specific genes provide better performance in downstream tasks. We stress that this should be adapted as the community moves forward with ST data analysis. Please refer to supplementary for further details and discussion.
When evaluating on gene prediction task, it is important to avoid sparse genes. We evaluated on both top 250 highly expressed genes overall (as in [17]) and top 250 highly expressed genes within our cancer-specific gene lists.
4.1 Quantitative Results
Figure 5: Comparison of gene-image distances from 100 randomly sampled spot pairs. Each point represents the gene and image distance between two spots. A higher $R^{2}$ [52] indicates a stronger linear correlation, suggesting better alignment between gene and image features.
Figure 6: Visualization of FASN gene expression predictions from different methods, with all values normalized to the range of 0 to 1.
Figure 7: Rank accuracy for different methods during training.
Gene Expression Prediction. As shown in Table 1, RankByGene outperforms all baseline methods across the Breast-ST1, Breast-ST2, and Lung-ST datasets. Specifically, RankByGene achieves a PCC improvement of $9\%$ to $34\%$ for the top 250 highly expressed genes and $7\%$ to $35\%$ for the top 250 cancer-specific genes compared to the best baseline performance in each dataset. We also provide the results evaluated on the full cancer-specific gene list in the supplementary. These results indicate that our method achieves a stronger alignment between image and gene expression data, with the learned gene-informed image features effectively capturing the underlying global relationships between spots.
Gene-related Classification. Using models trained on Breast ST and Lung ST datasets, we evaluate their performance on BCNB and TCGA-LUAD mutation datasets, respectively, as shown in Figure 4(a). Each method’s image encoder extracts patch-level features, which are then aggregated using the ABMIL [30] for slide-level prediction. RankByGene achieves strong AUC performance on most mutation statuses, demonstrating that enhanced image-gene alignment enables the model to capture gene-related information more effectively. This highlights the advantage of integrating gene information for downstream tasks.
Survival Analysis. Following the same setup, we evaluate survival prediction on TCGA-BRCA and TCGA-LUAD WSIs. Figure 4(b) shows that RankByGene achieves the highest C-Index (0.6814 for TCGA-BRCA and 0.5945 for TCGA-LUAD), further demonstrating its ability to capture survival-related signals and potential for complex downstream tasks.
4.2 Ablation Study
In our ablation studies, we systematically assess the influence of each key component and the selected hyperparameters. First, we vary the weights of the cross-modal ranking consistency loss and the intra-modal distillation loss to analyze their individual effects on performance. We also perform a comprehensive ablation study on the inclusion of each loss component, illustrating the contribution of each module to the overall alignment quality. Furthermore, we compare different image encoders, including CNN-based and Transformer-based to verify the generalizability of our method. The results are shown in Table 3. A detailed examination is provided in the supplementary.
@@ 4.3 Qualitative Result
Rank Accuracy During Training. We define Rank Accuracy as a metric to evaluate whether the ranking relationships among gene features are effectively captured by the image features during training. After the end of each epoch, Rank Accuracy is calculated as follows: the first gene-image pair in each batch is randomly selected as the anchor pair, along with two additional target pairs. Rank Accuracy is determined by checking whether the target pair with higher gene feature similarity to the anchor pair also has a higher image feature similarity. This process is repeated eight times to calculate the final Rank Accuracy. In Figure 7, we compare the change of the rank accuracy during training between different methods. Additionally, in Figure 5, we randomly select gene-image pairs to examine whether the distance relationships in gene expression values are reflected in image feature distances. The two figures together demonstrate that our method achieves a larger improvement in accuracy during training compared to the baseline methods, suggesting that the rank relationships in the gene features are effectively captured by image features.
Visualization of Cancer Marker Gene. To evaluate the accuracy of our model’s predictions for critical genes, we selected FASN [24] as the marker gene for visualization on the Breast-ST1 dataset due to its association with poor prognosis in breast invasive carcinoma across two independent cohorts [4]. We visualized the gene prediction values for the FASN from different methods across each spot, including ground truth values and the corresponding WSI region, as shown in Figure 6. Compared to the baseline method, our approach provides more accurate predictions for the FASN, demonstrating its effectiveness in capturing important gene expression patterns. We also provide visualizations of other marker genes for breast and lung cancer in the supplementary materials.
@@ 5 Conclusion
We present a robust and scalable approach for aligning spatial transcriptomics data with histopathology images, addressing major obstacles in cross-modal alignment. This method employs a ranking-based mechanism to capture similarity relationships between gene and image features, ensuring resilience to distortions and effectiveness across local and global scales while preserving crucial spatial and molecular patterns. A self-supervised knowledge distillation procedure further refines alignment by handling high dimensionality, sparsity, and noise in gene expression data. Through extensive experiments, our framework not only improves alignment quality, but also enhances predictive performance in gene expression prediction, slide-level classification, and survival analysis. This work thus provides a foundational tool for advancing multi-modal representation learning in digital pathology.